OpenStudy (anonymous):

Midpoint of 0.5 and 2.5

OpenStudy (anonymous):

add them then divide by 2

OpenStudy (anonymous):

the midpoint would be 1.5 you would take 0.5 +2.5= 3/2 =1.5